Guthrie Enterprises needs someone to supply it with 142,000 cartons of machine screws per year to support its manufacturing needs over the next five years. It will cost $1,820,000 to install the equipment necessary to start production; you’ll depreciate this cost straight-line to zero over the project’s life. You estimate that in five years this equipment can be salvaged for $152,000. Your fixed production costs will be $267,000 per year, and your variable production costs should be $9.60 per carton. You also need an initial investment in net working capital of $132,000. The tax rate is 22 percent and you require a return of 12 percent on your investment. Assume that the price per carton is $16.20.

a. Calculate the project NPV.

b. What is the minimum number of cartons per year that can be supplied and still guarantee a zero NPV? Verify that the quantity you calculated is enough to at least have a zero NPV.

c. What is the highest fixed costs that could be incurred and still guarantee a zero NPV? Verify that the fixed costs you calculated are enough to at least have a zero NPV.
Answer #1

1.

=-1820000+152000*(1-22%)/1.12^5-132000+132000/1.12^5+((142000*(16.20-9.60)-267000-1820000/5)*(1-22%)+1820000/5)/12%*(1-1/1.12^5)

=363263.3397

2.
-1820000+152000*(1-22%)/1.12^5-132000+132000/1.12^5+((Q*(16.20-9.60)-267000-1820000/5)*(1-22%)+1820000/5)/12%*(1-1/1.12^5)=0

=>Q=(((1820000-152000*(1-22%)/1.12^5+132000-132000/1.12^5)*12%/(1-1/1.12^5)-1820000/5)/(1-22%)+1820000/5+267000)/(16.20-9.60)

=>Q=122424.8668

3.
-1820000+152000*(1-22%)/1.12^5-132000+132000/1.12^5+((142000*(16.20-9.60)-F-1820000/5)*(1-22%)+1820000/5)/12%*(1-1/1.12^5)=0

=>F=-(((1820000-152000*(1-22%)/1.12^5+132000-132000/1.12^5)*12%/(1-1/1.12^5)-1820000/5)/(1-22%)+1820000/5-142000*(16.20-9.60))

=>F=396195.8791